OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of x2 = 0.

OpenStudy (anonymous):

A no real roots B one real root C two distinct real roots D three distinct real roots

OpenStudy (anonymous):

x^2=0

OpenStudy (jdoe0001):

\(\bf x^2 = 0\implies {\color{blue}{ 1}}x^2{\color{red}{ +0}}x{\color{green}{ +0}}=0\qquad discriminant\to \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}\) http://www.regentsprep.org/Regents/math/algtrig/ATE3/discriminant.htm

OpenStudy (anonymous):

im so confused

OpenStudy (jdoe0001):

well what did you get for the discriminant?

OpenStudy (anonymous):

0

OpenStudy (jdoe0001):

well so ohh shoot... ahemmm the discrimoiant should only be \(\bf {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}\) not root, my bad so notice, when you get 0, on the site on the link, it has a guideline for when it's + or - or 0

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

so, 0 means only 1 real root :)

OpenStudy (anonymous):

since i got zero does that mean there's only one

OpenStudy (anonymous):

ohhh :P

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

can i ask u another?? :)